A^{3}: A SIMPLE AND AYSMPTOTICALLY ACCURATE MODEL FOR COARSE-GRAINED MACHINES

Sanjay Ranka University of Florida

http://www.cis.ufl.edu/ ranka

March 1996 Outline of the talk

1. Important Features of coarse grained machines.

2. Classification of known models based on these features.

3. A^{3} (Approximate Model for Analysis of Aggregate Communication Operations) and Computation Model).

4. Performance of the model on key applications.

5. Limitations of the Model.